OpenStudy (anonymous):

cylinder a has a radius of 1 m and a height of 4 m. Cylinder B has a radius of 1 m and a height of 8m.What is the ratio of the volume of cylinder A to the Volume of cylinder B

OpenStudy (anonymous):

help plzz anyone

OpenStudy (anonymous):

Because \(V = \pi r^2 h\), you can write ratio like that: \(V_A:V_B = \pi r_A^2h_A:\pi r_B^2h_B = \boxed{r_A^2h_A:r_B^2h_B}\) \(r_A\) is radius of cylinder A \(h_A\) is height of cylinder A \(r_B\) is radius of cylinder B \(h_B\) is height of cylinder B Just plug in values and simplify

OpenStudy (anonymous):

these are my choices a.1:2 b.2:1 c.1:1 d.1:4